# Bridging the classical and quantum regimes in a dissipative Ising chain

### v1

@@ Abstract
We study the long-time dynamics of a dissipative Ising chain with varying quantum correlation. Invoking an ensemble-average formalism, and assuming spatial translation symmetry, we show that the dynamics can be described by a Lindblad master equation with an interpolated coherent Hamiltonian. In the classical limit, the interpolation Hamiltonian leads to a set of nonlinear equations of motion, where limit cycles can emerge in the long-time dynamics. In the quantum limit, by contrast, the system approaches a ferromagnetic steady state at long times. In between the two extremes, the discrete spatial translation symmetry can be spontaneously broken, as an antiferromagnetic steady state emerges, bridging the classical and quantum regimes. In particular, we illustrate how the classical limit-cycle behavior gradually disappears with the increase of quantum correlation. Since our model in the two extremes respectively applies to a dissipative Rydberg gas in the high- and zero-temperature limits, we expect it to provide an effective description of dissipative Rydberg gases at interim temperatures, and shed light on the fate of limit cycles in a quantum open system.
@@ II Model
$\dot{\rho}=i[H,\rho]+\gamma\sum_{j}\left(L_{j}\rho L_{j}^{\dagger}-\frac{1}{2}\{L_{j}^{\dagger}L_{j},\rho\}\right),$ (1)
@@
Physically, the dissipative Ising chain above can be realized in a dissipative Rydberg-atom array [12, 13, 14, 15, 50, 4, 51], where the ground and Rydberg states of the $j$th atom form the basis of the $j$th spin. It follows that, $\Omega$ is the effective Rabi frequency of the coupling between the ground and Rydberg states, $\Delta$ is the detuning of the coupling, and $n_{j}$ corresponds to the number operator for the Rydberg state of the $j$th atom. The interaction derives from the Van der Waals interaction between adjacent Rydberg atoms.
@@
For this purpose, we consider an ensemble of systems with different cluster partitions. For a one-dimensional system of $N$ spins with nearest-neighbor interactions, there are $2^{N}$ possible partitions. Each partition can be uniquely labeled by a set $C_{j}=\{j_{1},j_{2},\ldots\}$, where the element $j_{i}$ represents the bond on which the local interaction $h_{j_{i},j_{i}+1}$ is replaced by its mean-field decoupled counterpart $h_{j_{i},j_{i}+1}^{\text{mf}}$. Here
@@
$\dot{\rho}=i[H(C_{j}),\rho]+\gamma\sum_{i}\left(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}\right):=\mathcal{L}_{j}\rho,$ (7)
@@
$=$ $i\sum_{j}P(C_{j})[H(C_{j}),\rho]+\gamma\sum_{i}\left(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}\right).$ (8)
@@
As the system is tuned from the classical limit toward the quantum limit, $\lambda$ increases smoothly from 00 to $1$. Hence $\lambda$ can be regarded as a measure of quantum correlation. Intuitively, $\lambda$ can be a monotonic function in temperature in a dissipative Rydberg gas.
@@ III Dynamic phase diagram
Figure 1: Schematic illustration and typical dynamics of the dynamical phases. (a)(c)(e) Schematic illustration of the limit cycle phase (LC), the antiferromagnetic phase (AFM), and the ferromagnetic phase (FM), respectively. Colors of the symbols indicate different spin-up occupations at long times. (b)(d)(f) Numerically simulated time evolution of $\rho_{\text{odd}}^{r}$ (orange line) and $\rho_{\text{even}}^{r}$ (blue line) under the parameters $\lambda=0.1,0.2,$ and $0.5$, which correspond to the LC, AFM, and FM phases, respectively. Other parameters are $\Omega=1.5$, $\Delta=2$, $V=5$, and $N=6$.
@@
For numerical simulations, we define the spin-up occupation on the $j$th site as $\rho_{j}^{r}=\langle n_{j}\rangle$, where the expectation value is taken with respect to the density matrix $\rho(t)$. We then calculate the average spin-up occupation for odd lattice sites, $\rho^{r}_{\text{odd}}=\frac{2}{N}\sum_{i=1}^{N/2}\rho^{r}_{2i-1}$, and even lattice sites, $\rho^{r}_{\text{even}}=\frac{2}{N}\sum_{i=1}^{N/2}\rho^{r}_{2i}$, respectively. Figure 1(b), (d), and (f) show the time evolution of $\rho_{\text{odd}}^{r}$ and $\rho_{\text{even}}^{r}$ for $\lambda=0.1,0.2,$ and $0.5$, respectively. These numerical results are obtained for a chain of $N=6$ atoms with a periodic boundary condition. For $\lambda\lesssim 0.17$, the system is in the LC phase, marked by sustained oscillations of $\rho_{j}^{r}$ over time, typical of a continuous time crystal. When $\lambda$ is in the range $0.17\leq\lambda\leq 0.43$, the system is in the AFM phase. For $\lambda\gtrsim 0.43$, the system enters the FM phase, where all atoms reach the same steady-state value of $\rho_{j}^{r}$.
Figure 2: (a)(b) Numerically evaluated $D$ and its directional derivative $\partial_{\mathbf{v}}D$ along the direction $\mathbf{v}=(1,0)$, i.e., increasing $\lambda$, as functions of $\lambda$. The red vertical dashed lines indicate the transition points between different dynamic phases. For the calculations here, we fix $\Delta=2$ and $V=5$. (c) The directional derivative $\partial_{\mathbf{v}}D$ with $\mathbf{v}=\frac{1}{\sqrt{2}}(-1,1)$, corresponding to increasing $\Delta$ and decreasing $\lambda$, plotted as a function of both $\lambda$ and $\Delta$ for $V=5$. The white dashed lines denote the phase boundaries, determined by discontinuities in $\partial_{\mathbf{v}}D$, as illustrated (b). (d) Phase diagrams in the $(\lambda,\Delta)$ plane for different values of $V$, where the phase boundaries are again identified by discontinuities in $\partial_{\mathbf{v}}D$. The dark blue region represents the FM phase, the cyan region corresponds to the AFM phase, and the red region indicates the LC phase. From left to right, the panels show the results for $V=1.5,3,4.5,$ and $6$, respectively. Other parameters are fixed at $\Omega=1.5$ and $N=6$.
@@
$D=\mathrm{Var}\left(\{\rho_{j}^{r}(t)|j=1,2,..N,t\in\{t_{i}\}\}\right),$ (12)
@@
In Fig. 2(a)(b), we show the function $D$ and its directional derivative $\partial_{\mathbf{v}}D$ along the direction $\mathbf{v}=(1,0)$, corresponding to increasing $\lambda$. The vector $\mathbf{v}$ is defined in the two-dimensional parameter space $(\lambda,\Delta)$, where the horizontal axis represents $\lambda$ and the vertical axis represents $\Delta$. While $D$ varies continuously across the phase boundaries, its directional derivative $\partial_{\mathbf{v}}D$ exhibits clear discontinuities at the boundaries (highlighted by red dashed lines). This is more visible in Fig. 2(c), where we compute the directional derivative along $\mathbf{v}=\frac{1}{\sqrt{2}}(-1,1)$, which corresponds to increasing $\Delta$ and decreasing $\lambda$. In this case, $\partial_{\mathbf{v}}D=\frac{1}{\sqrt{2}}(\partial_{\Delta}-\partial_{\lambda})D$. Distinct phase regions are revealed in the color contour and are separated by discontinuities in $\partial_{\mathbf{v}}D$ (white dashed lines).
We then show the phase diagrams in the ($\lambda,\Delta$) plane for different values of $V$ in Fig. 2(d). For $V=1.5$ (leftmost panel), the entire parameter space is occupied by the FM phase (dark blue). When $V=3$ (second panel to the left), the AFM phase emerges (cyan) at lower $\lambda$ values. As $V$ increases further (the rightmost two panels), a third region, the LC phase (red) appears at small $\lambda$, particularly around $\Delta\sim 2$. With increasing $V$, both phase boundaries shift toward larger $\lambda$ values, demonstrating the effect of interaction on the dynamic phase transitions. These are consistent with the understanding that small $V$ or large $\lambda$ tends to suppress the classical nonlinear effects.
@@ IV Steady-state phase transition
[…]
Figure 3: Steady-state correlation functions and entanglement negativity. (a) Variation of the steady-state correlation function $C_{r=3}$ as a function of system parameters, where the color represents $C_{3}$, and the dashed line denotes the AFM–FM phase boundary. (b) Log-log plot of the correlation function $C_{r}$ as a function of $r$, where different colored curves correspond to markers of the corresponding color in (a). (c) Steady-state entanglement negativity $\mathcal{N}(\rho_{s})$ as a function of system parameters, with the dashed line indicating the AFM–FM phase boundary. (d) Variation of entanglement negativity with $V/V_{c}$ for $\lambda=0.4,0.45,0.5$, where $V_{c}$ represents the AM–FM transition point. Different curves correspond to distinct values of $\lambda$, as indicated by the arrows of matching colors in (c). We set $N=6$ in (a)(c)(d), and $N=8$ in (b). The remaining parameters are fixed as $\Omega=1.5$, $\Delta=2$, with periodic boundary conditions.
In Fig. 3(a), we numerically evaluate the steady-state correlation function $C_{r=3}$ with periodic boundary condition for $N=6$. Apparently, the correlation function peaks near the AFM-FM phase boundary (white dashed). To further explore this behavior, we show a log-log plot of the correlation $C_{r}$ as a function of $r$ for a system size of $N=8$, as shown in Fig. 3(b). The dashed line with red square markers corresponds to parameters at the phase boundary, while the other two curves represent parameters on either side of the phase boundary [see corresponding markers Fig. 3(a)]. Notably, at the phase transition, the correlation function shows the slowest decrease with distance, indicating critical behaviors. However, when the phase boundary is well within the classical regime with small $\lambda$, the correlation function remains nearly zero and does not exhibit significant changes across the phase transition. This is expected, since for small $\lambda$, the coherent Hamiltonian is dominated by its mean-field component, leading to the relation $\langle\sigma_{i}^{z}\sigma_{i+r}^{z}\rangle\approx\langle\sigma_{i}^{z}\rangle\langle\sigma_{i+r}^{z}\rangle$.
In Fig. 3(c), we calculate the entanglement negativity of the steady state $\rho_{s}$, which is defined as
$\mathcal{N}(\rho_{s})=\sum_{\alpha=\text{odd},\text{even}}\left(\sum_{\lambda^{\alpha}_{i}<0}\lambda^{\alpha}_{i}\right),$ (14)
@@ V Frequency Bifurcation
Spectral analysis is one of the most common methods used to investigate oscillatory behavior in nonlinear systems [49]. In the previous sections, we observed that, for small values of $\lambda$, the long-time dynamics exhibit oscillatory behavior characteristic of limit cycles. Here we study the frequency spectrum $\rho(f)$ of the spin-up occupation, obtained by performing a fast Fourier transform (FFT) on the time series of $\rho_{\text{even}}^{r}(t)$. The FFT is applied to the time series taken within a time window following a sufficiently long evolution. This yields the frequency distribution $\rho(f)$, where $f$ denotes frequency.
@@
Figure 4: Frequency spectrum of the long-time density matrix, with $\Delta=2$, $\Omega=1.5$, and $N=6$. (a) The case of $V=5$, where only the LC and AFM phases are present. The vertical dashed line marks the Hopf bifurcation separating the two phases. (b) The case of $V=9$, where rich nonlinear behaviors emerge, as four distinct phases can be identitifed: the near-chaotic phase (nC), the period-doubled limit cycle phase (LC2), the single-period limit cycle phase (LC1), and the AFM phase. The dashed lines indicate the boundaries between these phases. Only the positive-frequency components are shown, and the color represents the logarithmic amplitude $\log[\rho(f)]$.
@@
In Fig. 2(c)(d), the transition between the LC and AFM phases is characterized by discontinuities in the derivative of $D$. An alternative and more standard identifier is through spectral analysis. In Fig. 4(a), we compute the logarithmic amplitude $\log[\rho(f)]$ as a function of frequency $f$ (only the positive-frequency components are shown) and the interpolation parameter $\lambda$, with $V=5$ and $\Delta=2$. We observe that for $\lambda\lesssim 0.17$ (vertical dashed line), there exists a dominant frequency component, a prominent signature of oscillations in the LC phase. For $\lambda>0.17$, this frequency component disappears, signifying the system approaching a steady state at long times. This transition corresponds to a Hopf bifurcation, where the limit-cycle oscillations vanish and the system settles into a stable fixed point [49].
@@
As we increase $V$ further, the frequency spectrum begins to show more complex nonlinear behaviors. In Fig. 4(b), we show $\log[\rho(f)]$ with a larger interaction strength $V=9$. For small $\lambda$, in the region labeled “nC”, the spectrum shows a wide, almost continuous range, typical of near-chaotic motion. This spectral broadening indicates that the system no longer exhibits well-defined periodic dynamics. As $\lambda$ increases, the distribution of the frequency components becomes sparser. Around $\lambda=0.04$, the spectrum is composed of the main frequency, and a frequency at half of the main one, indicating that the system is entering the LC phase. We label this region as “LC2”. As $\lambda$ increases and crosses the vertical dashed line near $\lambda\sim 0.08$, a period-doubling bifurcation occurs, the half-frequency components disappear, and the oscillation period is reduced by half. This marks a transition into the “LC1” region, where the system shows single-frequency limit-cycle dynamics. Finally, at $\lambda\sim 0.25$, a Hopf bifurcation occurs, where oscillatory dynamics disappear and the system enters the AFM state. Here what remains from the LC phase is only the spontaneously broken translational symmetry [see also Fig. 1(b)(d)]

### v2

@@ Abstract
We study the long-time dynamics of a dissipative Ising chain with varying quantum correlation. Invoking an ensemble-average formalism, and assuming spatial translation symmetry, we show that the dynamics can be described by a Lindblad master equation with an interpolated coherent Hamiltonian. In the classical limit, the interpolation Hamiltonian leads to a set of nonlinear equations of motion, where limit cycles can emerge in the long-time dynamics. In the quantum limit, by contrast, the system approaches a ferromagnetic steady state at long times. In between the two extremes, the discrete spatial translation symmetry can be spontaneously broken, as an antiferromagnetic steady state emerges, bridging the classical and quantum regimes. In particular, we illustrate how the classical limit-cycle behavior gradually disappears with the increase of quantum correlation. Since our model in the two extremes respectively applies to a dissipative Rydberg gas in the high- and zero-temperature limits, we expect it to provide a qualitatively correct description of dissipative Rydberg gases at interim temperatures, and shed light on the fate of limit cycles in a quantum open system.
@@ II Model
$\dot{\rho}=-i[H,\rho]+\gamma\sum_{j}\left(L_{j}\rho L_{j}^{\dagger}-\frac{1}{2}\{L_{j}^{\dagger}L_{j},\rho\}\right),$ (1)
@@
Physically, the dissipative Ising chain above can be realized in a dissipative Rydberg-atom array [12, 13, 14, 15, 50, 4, 51], where the ground and Rydberg state of the $j$th atom form the basis of the $j$th spin. The quantum-jump processes in Eq. (1) correspond to the spontaneous decay of the Rydberg states, which satisfy the Born-Markov approximation [52], justifying the application of the Lindblad master equation. Indeed, such a treatment is widely used for describing dissipative Rydberg atoms [24, 23, 25, 26, 28, 27, 31, 29, 30]. It also follows that, $\Omega$ is the effective Rabi frequency of the coupling between the ground and Rydberg states, $\Delta$ is the detuning of the coupling, and $n_{j}$ corresponds to the number operator for the Rydberg state of the $j$th atom. The interaction derives from the Van der Waals interaction between adjacent Rydberg atoms.
@@
For this purpose, we consider an ensemble of systems with different cluster partitions. For a one-dimensional system of $N$ spins with nearest-neighbor interactions under the periodic boundary condition, there are a total of $N$ local interaction terms $h_{i,i+1}$, and each can be switched on or off, leading to $2^{N}$ possible partitions. Each partition can be uniquely labeled by a set $C_{j}=\{j_{1},j_{2},\ldots\}$, where the element $j_{i}$ represents the bond on which the local interaction $h_{j_{i},j_{i}+1}$ is replaced by its mean-field decoupled counterpart $h_{j_{i},j_{i}+1}^{\text{MF}}$. Here
@@
$\dot{\rho}=-i[H(C_{j}),\rho]+\gamma\sum_{i}\left(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}\right):=\mathcal{L}_{j}\rho,$ (7)
@@
$=$ $-i\sum_{j}P(C_{j})[H(C_{j}),\rho]+\gamma\sum_{i}\left(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}\right).$ (8)
@@
Note that for the derivation above, the periodic boundary condition is imposed. As the system is tuned from the classical limit toward the quantum limit, $\lambda$ increases smoothly from 0 to $1$. Hence $\lambda$ can be regarded as a measure of quantum correlation. In particular, in the limit $\lambda=0$, the model is equivalent to substituting the product state $\rho=\bigotimes_{j}\rho_{j}$ into Eq. (1), under which there are no entanglements between spins on different sites [41]. Note that we have not systematically introduced temperature nor considered the thermal effects. But, since our model in the two limits $\lambda=0$ and $\lambda=1$ respectively recovers the widely used descriptions for dissipative Rydberg atoms at high and zero temperatures [22, 23, 53, 54], and since quantum correlations become increasingly suppressed at higher temperatures, we expect our calculations in the intermediate regime ($0<\lambda<1$) should provide a qualitatively correct description for dissipative Rydberg gases at finite temperatures. A full-fledged finite-temperature theory is desirable but beyond the scope of current work.
@@ III Dynamic phase diagram
Figure 1: Schematic illustration and typical dynamics of the dynamical phases. (a)(c)(e) Schematic illustration of the limit cycle phase (LC), the antiferromagnetic phase (AFM), and the ferromagnetic phase (FM), respectively. Colors of the symbols indicate different spin-up occupations at long times. (b)(d)(f) Numerically simulated time evolution of $n_{\text{odd}}^{r}$ (orange line) and $n_{\text{even}}^{r}$ (blue line) under the parameters $\lambda=0.1,0.2,$ and $0.5$, which correspond to the LC, AFM, and FM phases, respectively. Other parameters are $\Omega=1.5$, $\Delta=2$, $V=5$, and $N=6$.
@@
For numerical simulations, we define the spin-up occupation on the $j$th site as $n_{j}^{r}=\langle n_{j}\rangle$, where the expectation value is taken with respect to the density matrix $\rho(t)$. We then calculate the average spin-up occupation for odd lattice sites, $n^{r}_{\text{odd}}=\frac{2}{N}\sum_{i=1}^{N/2}n^{r}_{2i-1}$, and even lattice sites, $n^{r}_{\text{even}}=\frac{2}{N}\sum_{i=1}^{N/2}n^{r}_{2i}$, respectively. Figure 1(b), (d), and (f) show the time evolution of $n_{\text{odd}}^{r}$ and $n_{\text{even}}^{r}$ for $\lambda=0.1,0.2,$ and $0.5$, respectively. These numerical results are obtained for a chain of $N=6$ atoms with a periodic boundary condition. For $\lambda\lesssim 0.17$, the system is in the LC phase, marked by sustained oscillations of $n_{j}^{r}$ over time, typical of a continuous time crystal. When $\lambda$ is in the range $0.17\leq\lambda\leq 0.43$, the system is in the AFM phase. For $\lambda\gtrsim 0.43$, the system enters the FM phase, where all atoms reach the same steady-state value of $n_{j}^{r}$.
Figure 2: (a)(b) Numerically evaluated $D$ and its directional derivative $\partial_{\mathbf{d}}D$ along the direction $\mathbf{d}=(1,0)$, which is relatively close to the normal directions of both phase boundaries at $\Delta=2$. The red vertical dashed lines indicate the transition points between different dynamic phases. For the calculations here, we fix $\Delta=2$ and $V=5$. (c) The directional derivative $\partial_{\mathbf{d}}D$ with $\mathbf{d}=\frac{1}{\sqrt{2}}(-1,1)$, plotted as a function of both $\lambda$ and $\Delta$ for $V=5$. Here $\mathbf{d}$ is chosen to align with the dominant normal direction of both phase boundaries. The white dashed lines denote the phase boundaries, determined by discontinuities in $\partial_{\mathbf{d}}D$, as illustrated (b). (d) Phase diagrams in the $(\lambda,\Delta)$ plane for different values of $V$, where the phase boundaries are again identified by discontinuities in $\partial_{\mathbf{d}}D$. The dark blue region represents the FM phase, the cyan region corresponds to the AFM phase, and the red region indicates the LC phase. From left to right, the panels show the results for $V=1.5,3,4.5,$ and $6$, respectively. Other parameters are fixed at $\Omega=1.5$ and $N=6$.
@@
$D=\mathrm{Var}\left(\{n_{j}^{r}(t)|j=1,2,..N,t\in\{t_{i}\}\}\right),$ (12)
@@
In Fig. 2(a)(b), we show the function $D$ and its directional derivative $\partial_{\mathbf{d}}D$ along the direction $\mathbf{d}=(1,0)$, corresponding to increasing $\lambda$. The vector $\mathbf{d}$ is defined in the two-dimensional parameter space $(\lambda,\Delta)$, where the horizontal axis represents $\lambda$ and the vertical axis represents $\Delta$. While $D$ varies continuously across the phase boundaries, its directional derivative $\partial_{\mathbf{d}}D$ exhibits clear discontinuities at the boundaries (highlighted by red dashed lines). This is more visible in Fig. 2(c), where we compute the directional derivative along $\mathbf{d}=\frac{1}{\sqrt{2}}(-1,1)$, which corresponds to increasing $\Delta$ and decreasing $\lambda$. In this case, $\partial_{\mathbf{d}}D=\frac{1}{\sqrt{2}}(\partial_{\Delta}-\partial_{\lambda})D$. Distinct phase regions are revealed in the color contour and are separated by discontinuities in $\partial_{\mathbf{d}}D$ (white dashed lines).
We then show the phase diagrams in the ($\lambda,\Delta$) plane for different values of $V$ in Fig. 2(d). For $V=1.5$ (leftmost panel), the entire parameter space is occupied by the FM phase (dark blue). When $V=3$ (second panel to the left), the AFM phase emerges (cyan) at lower $\lambda$ values. As $V$ increases further (the rightmost two panels), a third region, the LC phase (red) appears at small $\lambda$, particularly around $\Delta\sim 2$. With increasing $V$, both phase boundaries shift toward larger $\lambda$ values, demonstrating the effect of interaction on the dynamic phase transitions. These are consistent with the understanding that small $V$ or large $\lambda$ tends to suppress the nonlinear effects.
@@ IV Steady-state phase transition
where $\langle\cdot\rangle$ is the expectation value of steady state $\rho_{s}$.
Figure 3: Steady-state correlation functions and entanglement negativity. (a) Variation of the steady-state correlation function $C_{r=3}$ as a function of system parameters, where the color represents $C_{3}$, and the dashed line denotes the AFM–FM phase boundary. (b) Log-log plot of the correlation function $C_{r}$ as a function of $r$, where different colored curves correspond to markers of the corresponding color in (a). (c) Steady-state entanglement negativity $\mathcal{N}(\rho_{s})$ as a function of system parameters, with the dashed line indicating the AFM–FM phase boundary. (d) Variation of entanglement negativity with $V/V_{c}$ for $\lambda=0.4,0.45,0.5$, where $V_{c}$ represents the AFM–FM transition point. Different curves correspond to distinct values of $\lambda$, as indicated by the arrows of matching colors in (c). We set $N=6$ in (a)(c)(d), and $N=8$ in (b). The remaining parameters are fixed as $\Omega=1.5$, $\Delta=2$, with periodic boundary conditions.
In Fig. 3(a), we numerically evaluate the steady-state correlation function $C_{r=3}$ with periodic boundary condition for $N=6$. Apparently, the correlation function peaks near the AFM-FM phase boundary (white dashed). To further explore this behavior, we show a log-log plot of the correlation $C_{r}$ as a function of $r$ for a system size of $N=8$, as shown in Fig. 3(b). The dashed line with red square markers corresponds to parameters at the phase boundary, while the other two curves represent parameters on either side of the phase boundary [see corresponding markers Fig. 3(a)]. Notably, the correlation function at the phase boundary exhibits an approximately linear behavior in the log-log plot with distance $r$, consistent with an algebraic decay typical of quantum criticality. Furthermore, by increasing $\lambda$ across the phase boundary, the variation of $C_{r}$ for a given $r$ is not monotonic, peaking near the transition point. On the other hand, when the phase boundary is well within the classical regime with small $\lambda$, the correlation function remains nearly zero and does not exhibit significant changes across the phase transition. This is expected, since for small $\lambda$, the coherent Hamiltonian is dominated by its mean-field component, leading to the relation $\langle\sigma_{i}^{z}\sigma_{i+r}^{z}\rangle\approx\langle\sigma_{i}^{z}\rangle\langle\sigma_{i+r}^{z}\rangle$.
In Fig. 3(c), we calculate the entanglement negativity [57, 58, 59] of the steady state $\rho_{s}$, which is defined as
$\mathcal{N}(\rho_{s})=\sum_{\alpha=\text{odd},\text{even}}\left(\sum_{\varepsilon^{\alpha}_{i}<0}|\varepsilon^{\alpha}_{i}|\right),$ (14)
@@ V Frequency Bifurcation
Spectral analysis is one of the most common methods used to investigate oscillatory behavior in nonlinear systems [49]. In the previous sections, we observed that, for small values of $\lambda$, the long-time dynamics exhibit oscillatory behavior characteristic of limit cycles. Here we study the frequency spectrum $\tilde{n}(f)$ of the spin-up occupation, obtained by performing a fast Fourier transform (FFT) on the time series of $n_{\text{even}}^{r}(t)$. The FFT is applied to the time series taken within a time window following a sufficiently long evolution. This yields the frequency distribution $\tilde{n}(f)$, where $f$ denotes frequency.
@@
Figure 4: Frequency spectrum of the long-time density matrix, with $\Delta=2$, $\Omega=1.5$, and $N=6$. (a) The case of $V=5$, where only the LC and AFM phases are present. The vertical dashed line marks the Hopf bifurcation separating the two phases. (b) The case of $V=9$, where rich nonlinear behaviors emerge, as four distinct phases can be identitifed: the near-chaotic phase (nC), the period-doubled limit cycle phase (LC2), the single-period limit cycle phase (LC1), and the AFM phase. The dashed lines indicate the boundaries between these phases. Only the positive-frequency components are shown, and the color represents the logarithmic amplitude $\ln[\tilde{n}(f)]$.
@@
In Fig. 2(c)(d), the transition between the LC and AFM phases is characterized by discontinuities in the derivative of $D$. An alternative and more standard identifier is through spectral analysis. In Fig. 4(a), we compute the logarithmic amplitude $\ln[\tilde{n}(f)]$ as a function of frequency $f$ (only the positive-frequency components are shown) and the interpolation parameter $\lambda$, with $V=5$ and $\Delta=2$. We observe that for $\lambda\lesssim 0.17$ (vertical dashed line), there exists a dominant frequency component, a prominent signature of oscillations in the LC phase. For $\lambda>0.17$, this frequency component disappears, signifying the system approaching a steady state at long times. This transition corresponds to a Hopf bifurcation, where the limit-cycle oscillations vanish and the system settles into a stable fixed point [49].
@@
As we increase $V$ further, the frequency spectrum begins to show more complex nonlinear behaviors. In Fig. 4(b), we show $\ln[\tilde{n}(f)]$ with a larger interaction strength $V=9$. For small $\lambda$, in the region labeled “nC”, the spectrum shows a wide, almost continuous range, typical of near-chaotic motion. This spectral broadening indicates that the system no longer exhibits well-defined periodic dynamics. As $\lambda$ increases, the distribution of the frequency components becomes sparser. Around $\lambda=0.04$, the spectrum is composed of the main frequency, and a frequency at half of the main one, indicating that the system is entering the LC phase. We label this region as “LC2”. As $\lambda$ increases and crosses the vertical dashed line near $\lambda\sim 0.08$, a period-doubling bifurcation occurs, the half-frequency components disappear, and the oscillation period is reduced by half. This marks a transition into the “LC1” region, where the system shows single-frequency limit-cycle dynamics. Finally, at $\lambda\sim 0.25$, a Hopf bifurcation occurs, where oscillatory dynamics disappear and the system enters the AFM state. Here what remains from the LC phase is only the spontaneously broken translational symmetry [see also Fig. 1(b)(d)]